# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ III-B Constrained Arcs in Problem (1a)
Compared to unconstrained arcs, constrained arcs exhibit more complex behavior. In a constrained arc, $\exists 1\leq p\leq P$, s.t. ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$. The following proposition characterizes an unconstrained arc.
@@
If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in a constrained arc, then $\exists 1\leq r_{p}\leq n$, s.t. $\forall 1\leq r<r_{p}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$. During the constrained arc,
$u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},\,\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}},$ (10)
and
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\equiv 0.\end{dcases}$ (11)
Conversely, if the initial value ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{0}\right)+d_{p}=0,\\ \forall 1\leq r<r_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{0}\right)=0,\end{dcases}$ (12)
then under the driving of (10), ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds.
@@
Since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$, (2) implies the existence of the above $r_{p}$. Considering the $k$-th order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (11) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$. Hence, (10) holds.
@@
Assume that (10) holds and ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (12). Note that $\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}=0$; hence, the initial condition implies that (11) holds. ∎
In the above proposition, $r_{p}$ is called the order of the equality state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0$. Similarly, the analytical expression of a constrained arc can be obtained by Lemma 1.
Cases exist where multiple inequality constraints are active in a constrained arc. For this case, the active constraints should be consistent. Specifically, denote that $\mathcal{P}=\left\{p\in\left[P\right]:{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0\right\}$, where $\left[P\right]\triangleq\left\{1,2,\dots,P\right\}$. Then, $\forall p_{1},p_{2}\in\mathcal{P}$, $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{1}}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{1}}-1}{\boldsymbol{b}}}=\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{2}}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p_{2}}-1}{\boldsymbol{b}}}$ holds in the arc due to (10), and $\forall p\in\mathcal{P}$, (11) holds. The following example illustrates the above analysis:
@@
Consider an $n$-th order chain-of-integrators system, i.e.,
$\begin{dcases}\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right],\,\forall 1<k\leq n,\\ \dot{x}_{1}\left(t\right)=u\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right].\end{dcases}$ (13)
Two inequality constraints are introduced as $x_{1}+x_{3}\leq 1$ and $x_{2}\leq 0$. Assume that both $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ hold in a constrained arc. Then, $u\equiv-x_{2}=0$; hence, $x_{3}\equiv 1$, $x_{2}\equiv 0$, and $x_{1}\equiv 0$. Therefore, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent. Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are not consistent.
@@ III-C System Behavior
Following the terminology introduced in our previous work [8], the system behavior is defined as follows:
@@
A system behavior in problem (1a) is an arc with constant system dynamics, denoted as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$. During the arc, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$. $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{g}}}\equiv{\boldsymbol{0}}$ provides fully equality state constraints, where $\widehat{{\boldsymbol{F}}}$ has full row rank. $\mathcal{P}$ is the set of active inequality state constraints.
In Definition 1, an unconstrained arc can be denoted as ${\mathcal{S}}=\left({\boldsymbol{A}},u_{0}{\boldsymbol{b}},\sim,\sim,\varnothing\right)$, where $\sim$ means that no inequality constraints are introduced, and $u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ is a constant.
For a constrained arc, $\mathcal{P}\not=\varnothing$. (10) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{p}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}_{p}$, where
$\widehat{{\boldsymbol{A}}}_{p}={\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},\,\widehat{{\boldsymbol{b}}}_{p}={\boldsymbol{0}}.$ (14)
Although it allows that $\exists p_{1},p_{2}\in\mathcal{P}$, s.t. $\widehat{{\boldsymbol{A}}}_{p_{1}}\not=\widehat{{\boldsymbol{A}}}_{p_{2}}$, it is guaranteed by the consistency of active constraints that $\widehat{{\boldsymbol{A}}}_{p_{1}}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{A}}}_{p_{2}}{\boldsymbol{x}}$ during the arc. For convenience of discussion, let $\widehat{{\boldsymbol{A}}}$ be $\widehat{{\boldsymbol{A}}}_{p}$ with the minimum $p\in\mathcal{P}$, i.e., $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$. Evidently, $\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}}\right)$ is well-defined in Definition 1.
The active inequality state constraints state that $\forall p\in\mathcal{P}$, (11) holds. The above equalities induce $\left(\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$ where $\widehat{{\boldsymbol{F}}}$ has full row rank. For well-posedness, $\left(\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$ are determined by the equality constraints with minimum indices.
@@
Consider a chain-of-integrators system (13) with full box state constraints, i.e., $\forall 1\leq k\leq n$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $-x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. The $\left(2k-1\right)$-th constraint is $-x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$-th order, where $\forall 1\leq j<k$, $x_{j}\equiv 0$, and $u\equiv 0$. Then, $x_{k}\equiv x_{\mathrm{m}k}$ can be represented as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{I}}_{k\times n},x_{\mathrm{m}k}{\boldsymbol{e}}_{k},\left\{2k\right\}\right)$. Among them, $\widehat{{\boldsymbol{A}}}=\left(\hat{a}_{ij}\right)_{i,j=1}^{n}$, where $\hat{a}_{ij}=1$ if $k<i=j+1\leq n$, and otherwise, $\hat{a}_{ij}=0$.
In a system behavior ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ holds. Therefore, the analytical expression of ${\boldsymbol{x}}\left(t\right)$ in ${\mathcal{S}}$ can be given by the following lemma.
Lemma 1.
The initial value problem (IVP) where $\dot{\boldsymbol{x}}\left(t\right)={\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{b}}\left(t\right)$, $t\in{\mathbb{R}}$, ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ has a unique solution:
$\forall t\in{\mathbb{R}},\,{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}{\boldsymbol{b}}\left(\tau\right)\mathrm{d}\tau.$ (15)
Proof.
$\frac{\mathrm{d}}{\mathrm{d}t}\left({\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{x}}\left(t\right)\right)={\mathrm{e}}^{-{\boldsymbol{A}}t}\left(\dot{\boldsymbol{x}}\left(t\right)-{\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)\right)={\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{b}}\left(t\right)$ implies that ${\mathrm{e}}^{-{\boldsymbol{A}}t}{\boldsymbol{x}}\left(t\right)-{\mathrm{e}}^{-{\boldsymbol{A}}t_{0}}{\boldsymbol{x}}_{0}=\int_{t_{0}}^{t}{\mathrm{e}}^{-{\boldsymbol{A}}\tau}{\boldsymbol{b}}\left(\tau\right)\mathrm{d}\tau$. ∎
Lemma 1 implies that for an unconstrained arc where $u\equiv u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$,
For a constrained arc,
For a given system behavior, ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ has an explicit elementary expression since every scalar element of ${\boldsymbol{e}}^{\widehat{{\boldsymbol{A}}}t}$ is composed of exponentials and polynomials w.r.t. $t$.

### v2

@@ III-B Constrained Arcs in Problem (1a)
Compared to unconstrained arcs, constrained arcs exhibit more complex behavior. In a constrained arc, $\exists p\in\left[P\right]$, s.t. ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$. The following proposition characterizes a constrained arc.
@@
If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in a constrained arc for $t\in\left[t_{1},t_{2}\right]$, then $\exists r_{p}\in\left[n\right]$, s.t. $\forall r\in\left[r_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}{\boldsymbol{b}}=0$, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\not=0$. During the constrained arc, it holds that
$u=\widehat{{\boldsymbol{a}}}_{p}^{\top}{\boldsymbol{x}},\,\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}},\,\widehat{{\boldsymbol{a}}}_{p}\triangleq-\frac{\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\right)^{\top}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}},$ (10)
and $\forall t\in\left[t_{1},t_{2}\right]$,
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0,\\ \forall r\in\left[r_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}=0.\end{dcases}$ (11)
$r_{p}$ is called the order of the active constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$.
Conversely, if $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t. ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (11), then under the driving of (10), ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in $\left[t_{1},t_{2}\right]$.
@@
(2) implies ${\boldsymbol{c}}_{p}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\not={\boldsymbol{0}}$ since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$; hence, $r_{p}$ exists and serves as the minimum non-zero column index of a non-vzero matrix. For $r\in\left[r_{p}-1\right]$, considering the $r$th-order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (11) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$. Hence, (10) holds.
@@
Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t. ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (11). Under the driving of (10), note that $\frac{\mathrm{d}^{r_{p}}}{\mathrm{d}t^{r_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}=0$; hence, (11) holds in $\left[t_{1},t_{2}\right]$. ∎
It is allowed that multiple inequality constraints are active in a constrained arc, as proved in Proposition 3.
Lemma 1.
${\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$ and ${\boldsymbol{b}}\in{\mathbb{R}}^{n}$. The initial value problem where $\dot{\boldsymbol{x}}={\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}$, ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}$ has a unique solution:
$\forall t\in{\mathbb{R}},\,{\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\mathrm{d}\tau\right){\boldsymbol{b}}.$ (12)
Furthermore, if ${\boldsymbol{c}}\in{\mathbb{R}}^{n}$ satisfies $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$ holds, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$.
Proof.
The proof of (12) is omitted since it is elementary [31].
Assume that $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$. Considering the minimal polynomial of ${\boldsymbol{A}}$, $\exists p^{*}$, s.t. $p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$. Therefore, $\forall r\geq n$, $\exists p_{r}^{*}$ s.t. $p_{r}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{r}^{*}\left({\boldsymbol{A}}\right)$.
Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}\right)$. Since $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, Note that $\forall r\geq n+1$,
$\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)={\boldsymbol{c}}^{\top}p_{r-1}^{*}\left({\boldsymbol{A}}\right)\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)=0.$ (13)
It can be proved that ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)$ is analytic. By (13), $\forall t\in{\mathbb{R}}$,
${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{r=0}^{\infty}\frac{\left(t-t_{0}\right)^{r}}{r!}\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)\right|_{t_{0}}\equiv{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{0}\right).$ (14)
In other words, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$. ∎
Proposition 3.
[…]
$\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}=0,\\ \forall q\in\mathcal{P},\,r\in\left[n\right],\,{\boldsymbol{c}}_{q}^{\top}\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right)^{r}{\boldsymbol{x}}=0.\end{dcases}$ (15)
Active constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ for $p\in\mathcal{P}$ are called consistent if linear equations (15) have a feasible solution ${\boldsymbol{x}}$.
Conversely, assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t. ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (15). Then, under the driving of (10) with an arbitrary $p\in\mathcal{P}$, it holds that $\forall q\in\mathcal{P}$, ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}+d_{q}\equiv 0$ in $\left[t_{1},t_{2}\right]$.
Proof.
Assume that $\forall p\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in the constrianed arc. By Proposition 2, (10) holds $\forall p\in\mathcal{P}$. $\forall q\in\mathcal{P}$, since ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv{\mathrm{const}}$, (15) can be derived through considering the $r$th-order derivative of ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}$ for $r\in\left[n\right]$.
Assume that $\exists t_{0}\in\left[t_{1},t_{2}\right]$, s.t. ${\boldsymbol{x}}\left(t_{0}\right)$ satisfies (15). $\exists p\in\mathcal{P}$, (15) holds in $\left[t_{1},t_{2}\right]$. Then, $\forall q\in\mathcal{P}$, $r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$ holds. By Lemma 1, ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv-d_{q}$ holds. ∎
@@
Consider a chain-of-integrator system of order $n$, i.e.,
[…]
Two inequality constraints are introduced as $x_{1}+x_{3}\leq 1$ and $x_{2}\leq 0$. $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent, i.e., a constrained is allowed where both $x_{1}+x_{3}\equiv 0$ and $x_{2}\equiv 0$ hold. In this case, $x_{2}\equiv 0$ implies $x_{1}\equiv 0$ and $u\equiv 0$; hence, $x_{1}+x_{3}\leq 1$ implies $x_{3}\equiv 1$. Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are inconsistent, i.e., in a constrained arc, if $x_{1}+x_{3}\equiv 1$ holds, then $x_{2}\equiv 1$ cannot hold.
@@ III-C System Behavior
A system behavior in problem (1a) is an arc with constant system dynamics, denoted as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$. During the arc, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ holds. Induced by Proposition 3, ${\boldsymbol{F}}{\boldsymbol{x}}+{\boldsymbol{g}}\equiv{\boldsymbol{0}}$ provides full equality constraints, where ${\boldsymbol{F}}$ has full row rank. $\mathcal{P}$ is the set of consistent active state constraints.
In Definition 1, an unconstrained arc can be denoted as ${\mathcal{S}}=\left({\boldsymbol{A}},u_{0}{\boldsymbol{b}},\sim,\sim,\varnothing\right)$, where $\sim$ means that no equality constraints are introduced. $u_{0}\in\left\{\pm{u_{\mathrm{m}}}\right\}$ is a constant. By Lemma 1,
For a constrained arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$, $\mathcal{P}\not=\varnothing$. (10) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}$, where $\widehat{{\boldsymbol{a}}}_{p}$ is defined in (10). Although it allows that $\exists p,q\in\mathcal{P}$, s.t. $\widehat{{\boldsymbol{a}}}_{p}\not=\widehat{{\boldsymbol{a}}}_{q}$, it is guaranteed by the consistency of $\mathcal{P}$ that $\widehat{{\boldsymbol{a}}}_{p}^{\top}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{a}}}_{q}^{\top}{\boldsymbol{x}}$ during the arc. For convenience, let $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$. Then, ${\mathcal{S}}$ in Definition 1 is well-defined. By Lemma 1, during ${\mathcal{S}}$,
@@
Consider a chain-of-integrator system (16) with full box state constraints, i.e., $\forall k\in\left[n\right]$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $-x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. The $\left(2k-1\right)$-th constraint is $-x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$th-order, where $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$, and $u\equiv 0$. Then, $x_{k}\equiv x_{\mathrm{m}k}$ can be represented as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{I}}_{k\times n},x_{\mathrm{m}k}{\boldsymbol{e}}_{k},\left\{2k\right\}\right)$ where $\widehat{{\boldsymbol{A}}}=\left(\hat{a}_{ij}\right)_{i\in\left[n\right],j\in\left[n\right]}$. $\hat{a}_{ij}=1$ if $k<i=j+1\leq n$; otherwise, $\hat{a}_{ij}=0$.
Figure 1: A feasible BBS trajectory under the constraint ${\boldsymbol{x}}\geq{\boldsymbol{0}}$. The trajectory consists of 3 arcs, i.e., ${\mathcal{S}}_{1}$, ${\mathcal{S}}_{2}$, and ${\mathcal{S}}_{3}$. ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ is connected at ${\boldsymbol{0}}$, denoted as $\mathcal{E}_{1}$. The trajectory is tangent to $\left\{x_{3}=0\right\}$ at $\mathcal{T}_{1}^{(1)}$, $\mathcal{T}_{1}^{(2)}$, and $\mathcal{T}_{2}^{(2)}$.